what is the activity coefficient of each? use the table provided below:

What is the activity coefficient for each ion at the given ionic strength at 25 °C? Activity coefficients at various ionic strengths can be found in this table. = 0.05 M) Number 0.25 COA A0.001 M) Number 0.00022 Eu3+ ( 0.1 M) Number 0.022 (CH3CH2)3NH+ (μ = 0.05 M) Number Incorrect. 0.10
Answer #1

(a) CrO42- (u = 0.05 M) : Activity coefficient = 0.445

(b) Al3+ (u = 0.001 M) : Activity coefficient = 0.738

(c) Eu3+ (u = 0.1 M) : Activity coefficient = 0.18

(d) (CH3CH2)3NH+ (u = 0.05 M) : Activity coefficient = 0.83